OpenStudy (canadiansteve99):

solve the equation for x identify any extraneous solutions. √x = -3 A. There is no real solution B. -9 is a solution to the original equation. -9 is an extraneous solution. C. 9 is a solution to the original equation D. 9 is a solution to the original equation. -9 is an extraneous solution.

Nnesha (nnesha):

solve for x

OpenStudy (canadiansteve99):

How do I do that exactly?

Nnesha (nnesha):

how would you cancel out square root ?? do you konw ?

OpenStudy (canadiansteve99):

Hahaa no..sorry.

Nnesha (nnesha):

you should take square both side to cancel out square root and to cancel out square you should take square root

OpenStudy (canadiansteve99):

I really suck at math, so English please? Like step by step or something?

Nnesha (nnesha):

ohh i didn't notice i typed that in Japanese :( sorry about that you have to take square both side in order to cancel out square root

OpenStudy (canadiansteve99):

So square root of x and -3?

Nnesha (nnesha):

yes both side not the square root our goal is to solve for x right so we have to take square both side

Nnesha (nnesha):

remember you can convert root to exponent for example \[\huge\rm \sqrt{m} = m^\frac{ 1 }{ 2}\] take square both side \[\large\rm (m^\frac{ 1 }{ 2 })^2\] \[\large\rm m^\frac{ 2 }{ 2 }\] both 2 can cancel out so answer is m now do you understand why we have to take square ?

OpenStudy (canadiansteve99):

More so than I did earlier so thank you for that :D

OpenStudy (canadiansteve99):

Alright, I think I can figure this out now so, THANK YOU SO MUCH NNESHA!